OpenStudy (campayne):

Segment TU has endpoints T(0 , -3) and U(0, 1). Select the equation in point-slope form for the perpendicular bisector of TU. y - 1 = 0 y - 1= 2 y + 1 = -2 y + 1 = 0

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (campayne):

not in the slightest

OpenStudy (skullpatrol):

Do you know what the point-slope form looks like?

OpenStudy (campayne):

something like y- somethinng plus somethnig

OpenStudy (skullpatrol):

You need to memorize the exact form: $$\Huge y-y_0 =m(x-x_0)$$

OpenStudy (campayne):

but which numbers go where?

OpenStudy (skullpatrol):

The y sub 0 and x sub 0 are the coordinates of the point on the line and m is the slope of the line. That's why it's called the point-slope form of an equation of a line :-)

OpenStudy (campayne):

i see

OpenStudy (skullpatrol):

Also, note x and y are the variables, right?

OpenStudy (skullpatrol):

Now by drawing a sketch of T(0, -3) and U(0,1) you will see that the given line segment is on the y-axis and has length 4. Thus, the midpoint is M(0, -1). All you need now is the equation of a horizontal line that goes through that point because a horizontal line is perpendicular to a vertical line. The answer choice should be equivalent to y = -1. Look at y + 1 = 0 and subtract 1 from both sides.

OpenStudy (skullpatrol):

Looking back at the point-slope form and using m = 0, because all horizontal lines have a slope of 0 and the point M(0, -1) gives y + 1 = 0 as the answer in point-slope form.